# MIRAGE: A Benchmark for Multimodal Information-Seeking and Reasoning in Agricultural Expert-Guided Conversations

### v1

@@ E.5.6 Plant Care and Gardening Guidance Results (MMST Standard)
| gpt-4.1 | 3.38 | 3.73 | 3.55 | 3.01 | 3.45 | 3.73 | 3.17 | 3.12 | 3.52 | 3.80 | 3.43 | 3.24 |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | ||||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Each metric is scored on a 0–4 scale: Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple values indicate the best performance within each column block.
E.5.7 Weeds/Invasive Plants Management Results (MMST Standard)
@@
| gpt-4.1 | 2.90 | 3.41 | 3.14 | 3.01 | 3.37 | 2.87 | 2.89 | 3.40 | 3.00 | |||
| gpt-4.1-mini | ||||||||||||
| gpt-4o | 2.95 | 3.17 | 3.11 | |||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0–4 across Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple marks the highest score for each metric within a column block.
E.5.8 Plant Disease Management Results (MMST Contextual)
@@
| gpt-4.1 | 3.07 | 3.52 | 3.34 | 2.86 | 3.16 | 3.49 | 2.91 | 3.09 | 3.44 | 3.08 | 2.96 | |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | 3.17 | |||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0–4 across Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple marks the highest score for each metric within a column block.
E.5.9 Insect and Pest Management Results (MMST Contextual)
@@
| gpt-4.1 | 3.14 | 3.59 | 3.32 | 2.99 | 3.25 | 3.61 | 2.99 | 3.06 | 3.22 | 3.60 | 3.20 | 3.21 |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | ||||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple highlights the best score for each metric within a column block.
E.5.10 Plant Care and Gardening Guidance Results (MMST Contextual)
@@
| gpt-4.1 | 3.40 | 3.73 | 3.54 | 3.46 | 3.72 | 3.16 | 3.53 | 3.80 | 3.42 | 3.27 | ||
| gpt-4.1-mini | ||||||||||||
| gpt-4o | 3.08 | |||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | 3.50 | |||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple highlights the best score for each metric within a column block.
E.5.11 Weeds/Invasive Plants Management Results (MMST Contextual)
@@
| gpt-4.1 | 3.03 | 3.53 | 3.27 | 2.92 | 3.12 | 3.46 | 2.92 | 3.03 | 3.02 | 3.45 | 3.07 | 3.20 |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | ||||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple marks the highest score for each metric within a column block.
Appendix F MIRAGE-MMMT
F.1 Benchmark Details
@@ F.2.2 Output
F.2.3 Goal Inference and Decision Policy
@@ F.3 Evaluation Criteria
F.4 Data Curation Details
@@
Release Protocol for MIRAGE-MMMT: In designing our dataset release, we follow established best practices from recent benchmarks such as MMLU [35], and BIG-Bench [32, 33], which emphasize the importance of separating training data and test targets to prevent leakage and ensure reliable model evaluation. We adopt a protocol that maximizes transparency, reproducibility, and community usability, while preserving the integrity of the held-out test set. We publicly release:
@@
F.5 Additional MIRAGE-MMMT Results
@@
Appendix G Prompts
G.1 Evaluation Prompts for MIRAGE-MMST
@@
G.2 Evaluation Prompts for MIRAGE-MMMT
Appendix H Case Study
[…]
H.1 Category-Wise Cases
H.1.1 Plant Identification (MMST Standard)
H.1.2 Insect and Pest Identification (MMST Standard)
H.1.3 Plant Disease Identification (MMST Standard)
H.1.4 Plant Disease Management (MMST Standard)
H.1.5 Insect and Pest Management (MMST Standard)
H.1.6 Plant Care and Gardening Guidance (MMST Standard)
H.1.7 Weeds/Invasive Plants Management (MMST Standard)
H.1.8 Plant Disease Management (MMST Contextual)
H.1.9 Insect and Pest Management (MMST Contextual)
H.1.10 Plant Care and Gardening Guidance (MMST Contextual)
H.1.11 Weeds/Invasive Plants Management (MMST Contextual)
H.1.12 Clarify (MMMT)
H.1.13 Respond (MMMT)
H.2 Examples of Reasoning LLM as a Judge
H.2.1 MMST Identification Task
H.2.2 MMST Management Task
H.2.3 MMMT Task

### v2

@@ F.5.6 Plant Care and Gardening Guidance Results (MMST Standard)
| gpt‑4.1 | 3.38 | 3.73 | 3.55 | 3.01 | 3.45 | 3.73 | 3.17 | 3.12 | 3.52 | 3.80 | 3.43 | 3.24 |
| gpt‑4.1‑mini | ||||||||||||
| gpt‑4o | ||||||||||||
| gpt‑4o‑mini | ||||||||||||
| nova‑pro | ||||||||||||
| claude‑3‑7‑sonnet | ||||||||||||
| claude‑3‑5‑sonnet | ||||||||||||
| claude‑3‑haiku | ||||||||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
| gemma‑3‑4b‑it | ||||||||||||
| gemma‑3‑12b‑it | ||||||||||||
| gemma‑3‑27b‑it | ||||||||||||
| InternVL3‑2B | ||||||||||||
| InternVL3‑8B | ||||||||||||
| InternVL3‑14B | ||||||||||||
| InternVL3‑38B | ||||||||||||
| InternVL3‑78B | ||||||||||||
Models are color‑coded by type: closed‑source (red) and open‑source (blue). Each metric is scored on a 0–4 scale: Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple values indicate the best performance within each column block.
F.5.7 Weeds/Invasive Plants Management Results (MMST Standard)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
@@ F.5.7 Weeds/Invasive Plants Management Results (MMST Standard)
| gpt‑4.1 | 2.90 | 3.41 | 3.14 | 3.01 | 3.37 | 2.87 | 2.89 | 3.40 | 3.00 | |||
| gpt‑4.1‑mini | ||||||||||||
| gpt‑4o | 2.95 | 3.17 | 3.11 | |||||||||
| gpt‑4o‑mini | ||||||||||||
| nova‑pro | ||||||||||||
| claude‑3‑7‑sonnet | ||||||||||||
| claude‑3‑5‑sonnet | ||||||||||||
| claude‑3‑haiku | ||||||||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
| gemma‑3‑4b‑it | ||||||||||||
| gemma‑3‑12b‑it | ||||||||||||
| gemma‑3‑27b‑it | ||||||||||||
| InternVL3‑2B | ||||||||||||
| InternVL3‑8B | ||||||||||||
| InternVL3‑14B | ||||||||||||
| InternVL3‑38B | ||||||||||||
| InternVL3‑78B | ||||||||||||
Models are color‑coded by type: closed‑source (red) and open‑source (blue). Scores range from 0–4 across Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple marks the highest score for each metric within a column block.
F.5.8 Plant Disease Management Results (MMST Contextual)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
@@ F.5.8 Plant Disease Management Results (MMST Contextual)
| gpt‑4.1 | 3.07 | 3.52 | 3.34 | 2.86 | 3.16 | 3.49 | 2.91 | 3.09 | 3.44 | 3.08 | 2.96 | |
| gpt‑4.1‑mini | ||||||||||||
| gpt‑4o | 3.17 | |||||||||||
| gpt‑4o‑mini | ||||||||||||
| nova‑pro | ||||||||||||
| claude‑3‑7‑sonnet | ||||||||||||
| claude‑3‑5‑sonnet | ||||||||||||
| claude‑3‑haiku | ||||||||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
| gemma‑3‑4b‑it | ||||||||||||
| gemma‑3‑12b‑it | ||||||||||||
| gemma‑3‑27b‑it | ||||||||||||
| InternVL3‑2B | ||||||||||||
| InternVL3‑8B | ||||||||||||
| InternVL3‑14B | ||||||||||||
| InternVL3‑38B | ||||||||||||
| InternVL3‑78B | ||||||||||||
Models are color‑coded by type: closed‑source (red) and open‑source (blue). Scores range from 0–4 across Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple marks the highest score for each metric within a column block.
F.5.9 Insect and Pest Management Results (MMST Contextual)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
@@ F.5.9 Insect and Pest Management Results (MMST Contextual)
| gpt‑4.1 | 3.14 | 3.59 | 3.32 | 2.99 | 3.25 | 3.61 | 2.99 | 3.06 | 3.22 | 3.60 | 3.20 | 3.21 |
| gpt‑4.1‑mini | ||||||||||||
| gpt‑4o | ||||||||||||
| gpt‑4o‑mini | ||||||||||||
| nova‑pro | ||||||||||||
| claude‑3‑7‑sonnet | ||||||||||||
| claude‑3‑5‑sonnet | ||||||||||||
| claude‑3‑haiku | ||||||||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
| gemma‑3‑4b‑it | ||||||||||||
| gemma‑3‑12b‑it | ||||||||||||
| gemma‑3‑27b‑it | ||||||||||||
| InternVL3‑2B | ||||||||||||
| InternVL3‑8B | ||||||||||||
| InternVL3‑14B | ||||||||||||
| InternVL3‑38B | ||||||||||||
| InternVL3‑78B | ||||||||||||
Models are color‑coded by type: closed‑source (red) and open‑source (blue). Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple highlights the best score for each metric within a column block.
F.5.10 Plant Care and Gardening Guidance Results (MMST Contextual)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
@@ F.5.10 Plant Care and Gardening Guidance Results (MMST Contextual)
| gpt‑4.1 | 3.40 | 3.73 | 3.54 | 3.46 | 3.72 | 3.16 | 3.53 | 3.80 | 3.42 | 3.27 | ||
| gpt‑4.1‑mini | ||||||||||||
| gpt‑4o | 3.08 | |||||||||||
| gpt‑4o‑mini | ||||||||||||
| nova‑pro | ||||||||||||
| claude‑3‑7‑sonnet | ||||||||||||
| claude‑3‑5‑sonnet | ||||||||||||
| claude‑3‑haiku | ||||||||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
| gemma‑3‑4b‑it | ||||||||||||
| gemma‑3‑12b‑it | ||||||||||||
| gemma‑3‑27b‑it | ||||||||||||
| InternVL3‑2B | ||||||||||||
| InternVL3‑8B | ||||||||||||
| InternVL3‑14B | ||||||||||||
| InternVL3‑38B | ||||||||||||
| InternVL3‑78B | 3.50 | |||||||||||
Models are color‑coded by type: closed‑source (red) and open‑source (blue). Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple highlights the best score for each metric within a column block.
F.5.11 Weeds/Invasive Plants Management Results (MMST Contextual)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
@@ F.5.11 Weeds/Invasive Plants Management Results (MMST Contextual)
| gpt‑4.1 | 3.03 | 3.53 | 3.27 | 2.92 | 3.12 | 3.46 | 2.92 | 3.03 | 3.02 | 3.45 | 3.07 | 3.20 |
| gpt‑4.1‑mini | ||||||||||||
| gpt‑4o | ||||||||||||
| gpt‑4o‑mini | ||||||||||||
| nova‑pro | ||||||||||||
| claude‑3‑7‑sonnet | ||||||||||||
| claude‑3‑5‑sonnet | ||||||||||||
| claude‑3‑haiku | ||||||||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
| gemma‑3‑4b‑it | ||||||||||||
| gemma‑3‑12b‑it | ||||||||||||
| gemma‑3‑27b‑it | ||||||||||||
| InternVL3‑2B | ||||||||||||
| InternVL3‑8B | ||||||||||||
| InternVL3‑14B | ||||||||||||
| InternVL3‑38B | ||||||||||||
| InternVL3‑78B | ||||||||||||
Models are color‑coded by type: closed‑source (red) and open‑source (blue). Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple marks the highest score for each metric within a column block.
Appendix G MIRAGE-MMMT
G.1 Benchmark Details
@@ G.2.2 Output
G.2.3 Goal Inference and Decision Policy
@@ G.3 Evaluation Criteria
G.4 Data Curation Details
@@
Release Protocol for MIRAGE-MMMT: In designing our dataset release, we follow established best practices from recent benchmarks such as MMLU [hendrycks2020measuring], and BIG-Bench [ghazal2013bigbench, ghazal2017bigbench], which emphasize the importance of separating training data and test targets to prevent leakage and ensure reliable model evaluation. We adopt a protocol that maximizes transparency, reproducibility, and community usability, while preserving the integrity of the held-out test set. We publicly release:
@@
G.5 Additional MIRAGE-MMMT Results
@@
Appendix H Prompts
H.1 Evaluation Prompts for MIRAGE-MMST
@@
H.2 Evaluation Prompts for MIRAGE-MMMT
Appendix I Case Study
I.1 Category-Wise Cases
I.1.1 Plant Identification (MMST Standard)
I.1.2 Insect and Pest Identification (MMST Standard)
I.1.3 Plant Disease Identification (MMST Standard)
I.1.4 Plant Disease Management (MMST Standard)
I.1.5 Insect and Pest Management (MMST Standard)
I.1.6 Plant Care and Gardening Guidance (MMST Standard)
I.1.7 Weeds/Invasive Plants Management (MMST Standard)
I.1.8 Plant Disease Management (MMST Contextual)
I.1.9 Insect and Pest Management (MMST Contextual)
I.1.10 Plant Care and Gardening Guidance (MMST Contextual)
I.1.11 Weeds/Invasive Plants Management (MMST Contextual)
I.1.12 Clarify (MMMT)
I.1.13 Respond (MMMT)
I.2 Examples of Reasoning LLM as a Judge
I.2.1 MMST Identification Task
I.2.2 MMST Management Task
I.2.3 MMMT Task